ange Frida 0 one bird nge one hird with fourteen fishPlease help!!!!! Need the answer ASAP
Answer #1

Opportunity cost of catching 1 bird (Crusoe) = 24/2 = 12 fish

Opportunity cost of catching 1 bird (Friday) = 6/1 = 6 fish

Friday offers to exchange one bird for eleven fish with Robinson Crusoe. (option A)

(Any exchange which falls between the opportunity costs of the two would benefit both)